OpenStudy (anonymous):

2 + (a + b) = (2 + a) + b is an example of which algebraic property? Distributive Property Associative Property of Addition Commutative Property of Addition Symmetric Property

OpenStudy (hoblos):

it is the Associative Property of Addition a+b+c = a+(b+c) = (a+b)+c

OpenStudy (anonymous):

when you think of associative, think grouping (the parenthesis) when you think of commutative, think order --> 2+3 = 3+2